# Moduli stacks of Higgs bundles on stable curves

### v2

@@ Abstract
In this article, we construct a flat degeneration of the derived moduli stack of Higgs bundles on smooth curves using the stack of expanded degenerations of Jun Li. We show an intrinsic relative log-symplectic form on the degeneration, and we compare it with the one constructed by the second author. We show that the Hitchin map of the degeneration we construct has complete fibers. Furthermore, we show that the Hitchin map is flat and that a suitable open subset of the smooth locus of the reduced nilpotent cone is Lagrangian. We also extend the construction of the moduli of Higgs bundles along with the relative log-symplectic form over the universal moduli stack of stable curves.
1 Introduction
@@
It is well known that moduli stacks of vector bundles on nodal curves are not complete. Often, their completions involve adding torsion-free sheaves as points of the boundary. Because we want to use the mapping space techniques of shifted symplectic geometry, we adopt the alternative approach of completing the moduli by adding boundary points, which are bundles on a bubbling node. This approach goes back to the classical works of David Gieseker [12] and Jun Li [17]. We will implement the approach in our setting by utilizing Jun Li’s stack of expanded degenerations. This allows us to use only vector bundles instead of torsion-free coherent sheaves and apply techniques of shifted symplectic geometry [32], obtaining a relative symplectic structure on a complete moduli stack. We extend the Hitchin map to Higgs bundles on the bubbled curves parametrized by the stack of expanded degenerations and then use the completeness of this moduli stack to prove that our extended Hitchin map is complete.
The derived moduli stack of Higgs bundles on a smooth curve $C$ has virtual dimension
$-\chi_{\text{Dol}}(C,\mathcal{E}nd(E,\phi))=n^{2}\deg(c_{1}(T_{C}))=2n^{2}(g-1)$
@@
where $g$ is the genus of $C$ while the derived moduli stack of vector bundles (or any other Lagrangian in the derived moduli stack of Higgs bundle) has virtual dimension
$-\chi(Y,\mathcal{E}nd(E))=\int_{C}\text{ch}(\mathcal{E}nd(E))\text{td}(C)=n^{2}(g-1).$
@@
These moduli stacks are homotopically locally of finite presentation, and so exhibit the “hidden smoothness” envisioned in derived deformation theory. Given this it is natural to study other geometric properties of these moduli. This article investigates relative 00-shifted symplectic structures on the total space of a degeneration of $C$ to a nodal curve.
@@
After writing this article, we realized there are a lot of relations with the articles [2], [3] and significant overlap with a forthcoming article [11] of Ron Donagi and Andres Fernandez Herrero. In their work, they give a construction of a good moduli space for the semistable locus and semistable reduction relative to the Hitchin fibration (so the moduli space is proper over the Hitchin base in families). This uses the “infinite dimensional GIT” picture developed with Dan Halpern-Leistner. They also showed the flatness of the Hitchin morphism and that things are syntonic (as classical stacks), including the symplectic leaves. They are pursuing the story with punctures, thinking about the log Poisson reduction picture for the relative log cotangent stack of framed Gieseker vector bundles.
@@
To start with, we fix a family of projective curves $\mathcal{X}$ over the spectrum of a discrete valuation ring $S$ such that the generic fibre is smooth, the closed fibre is an irreducible nodal curve with a single node, and the total space $\mathcal{X}$ is smooth. We also fix a rank $n$ and degree $d$ for the vector bundles in our moduli problem. In Section §2, we recall the construction of the stack $\mathfrak{M}$ of bounded expanded degenerations (bounded by the integer $n$) of the family of curves $\mathcal{X}/S$ following [17] and [42]. One of the main results in this subsection is Lemma 3.9, which we use to prove the following proposition.
@@
Proposition 3.11  The morphism $\mathfrak{M}\longrightarrow S$ of ordinary Artin stacks is a log-smooth map. Moreover, the relative log-cotangent complex $\mathbb{L}^{log}_{\mathfrak{M}/S}=0$.
In subsection §2.3, we discuss shifted log-symplectic structures on quasi-smooth derived Artin stacks (Definition 3.4.0(a)) equipped with a locally free log structure. In subsection §2.4, we recall the definition of relative shifted symplectic forms for a quasi-smooth morphism of derived Artin stacks. We define relative shifted log-symplectic forms for certain logarithmic morphisms of derived Artin stacks equipped with locally-free log structures.
@@
In Section §3, we construct a logarithmic version of the relative Dolbeault moduli stack for the universal expanded family of curves $\mathcal{\mathcal{missing}}X_{\mathfrak{M}}\rightarrow\mathfrak{M}$. We show that the relative logarithmic Dolbeault moduli stack has a relative 00-shifted log-symplectic form over $S$. Moreover, we show that the relative log-symplectic form is an extension of Hitchin’s symplectic form on the generic fibre of the moduli stack over $S$. This was proved for moduli schemes in [9]. The main results of this section are the following.
@@
Theorem 4.7  $\mathcal{X}_{Dol}$ is $\mathcal{O}$-compact and $\mathcal{O}$-oriented over $S$. Hence, $\mathsf{Map}_{\mathcal{S}}(\mathcal{X}_{Dol},BGL_{n}\times\mathcal{S})$ has a 00-shifted relative symplectic structure over $S$.
@@
Theorem 4.9  The derived Artin stack $\mathsf{Map}_{{}_{\mathfrak{M}}}(\mathcal{X}_{\mathfrak{M},Dol},BGL_{n}\times\mathfrak{M})$ has a natural relative 00-shifted log-symplectic structure over $S$.
@@
In Section §4, we define the Hitchin map on the classical Artin stack of Gieseker-Higgs bundles $M^{cl}_{Gie}$ (see 8.1.2). We prove that the Hitchin map is complete.
Theorem 5.3  The morphism $h:M^{cl}_{Gie}\longrightarrow B$ is complete.
@@
In section §5, we study the reduced global nilpotent cone of $M^{cl}_{Gie}$, which is the reduction of the scheme theoretic fibre over the point $0\in B$. We prove that every irreducible component of of the reduced nilpotent cone has an open subset (denoted by $\mathcal{N}ilp^{sm,gen}$), which is an isotropic substack of $M$ (the derived stack of Higgs bundles) with respect to its log-symplectic form. We use this to compute the dimension of the reduced nilpotent cone and to show that the Hitchin map is flat. The main theorem in this section is the following.
@@
Theorem 7.6  There is a 00-shifted relative log-symplectic form on $\mathcal{M}^{Dol}_{g}$ (relative to the moduli stack of stable curves $\overline{\mathcal{M}_{g}}$).
In the Appendix, we construct the relative classical Artin stack of Gieseker-Higgs bundles and study its local properties. The main results of the appendix are Proposition 8.10 and Theorem 8.11. In the first lemma, we prove that the stack of Gieseker vector bundles is an almost very good stack. We use this in Lemma 8.11 to show that the classical stack of Gieseker-Higgs bundles is an irreducible, local complete intersection.
Proposition 8.10  The closed fibre $N^{cl}_{Gie,0}$ is an irreducible, equidimensional, almost very good stack ([37, Definition 2.1.2]) with normal crossing singularities.
Theorem 8.11  The stack $M^{cl}_{Gie,0}$ is an irreducible local complete intersection of pure dimension $12\dim N_{Gie,0}+$ 1.
1.1 Notations and Conventions
@@ 6. $\bullet$
    $\mathbf{dg}$, $\mathrm{cdga}$ (respectively $\mathrm{cdga}^{\leq 0}$) are endowed with their natural model structures for which equivalences are quasi-isomorphisms, and fibrations are epimorphisms (respectively, epimorphisms in strictly negative degrees).
@@ 8. $\bullet$
    The $\infty$-categories associated with the model categories $\mathbf{dg},\mathrm{cdga}^{\leq 0},\operatorname{dAff}$ are denoted by $\textbf{dg},\textbf{cdga}^{\leq 0},\textbf{dAff}$.
@@ 10. $\bullet$
    The $\infty$-category of derived stacks over $\mathbb{k}$, for the étale topology, is denoted by dSt. If $X$ is a derived stack, the $\infty$-category of derived stacks over $X$ is denoted by $\textbf{dSt}_{X}$. The truncation of a derived stack $X$ is denoted by $t_{0}X$.
@@ 12. $\bullet$
    List notations for log structures and classifying stacks of log structures.
  13. $\bullet$
@@
    ${\mathfrak{X}_{Dol}}$${\mathfrak{X}}$${T}$$\scriptstyle{q}$$\scriptstyle{p}$$\scriptstyle{r}$
2 Acknowledgements
T.P. (University of Pennsylvania) was partially supported by NSF FRG grant DMS- 2244978, NSF/BSF grant DMS-2200914, NSF grant DMS-1901876, and Simons Collaboration grant number 347070. O.B. (University of Haifa) and S.D. would like to acknowledge the Binational Science Foundation and NSF/BSF grant 2021717 for supporting S.D. as a postdoc. S.D. would like to thank Professor Alek Vainshtein of the University of Haifa, Israel, for the financial support from his Israel Science Foundation grant number 876/20 during a postdoc position at the University of Haifa. S.D. would like to thank Professor Vikraman Balaji for the financial support from his SERB Core Research Grant CRG/2022/000051-G during a postdoc position at Chennai Mathematical Institute. S.D. would also like to thank IISER Tirupati for their support during a part of this work.
3 Preliminaries
3.1 Space of bounded Expanded degenerations
In this subsection, we will recall a construction by Jun Li [17] called the stack of expanded degenerations. We will use this stack to construct our degeneration. The construction of Jun Li starts with a one-parameter family of varieties (of any dimension) such that the total space of the family is smooth, the generic fibre is smooth, and the special fibre is a normal crossing divisor in the total space of the family. For our purpose, we only need to consider the expanded degenerations of such a family for curves. We will make use of the standard fact that for any nodal curve, one can always construct a smoothing over the spectrum of a discrete valuation ring, which also has a smooth total space. The precise setting we consider can be spelled out as follows.
3.1.1 Degeneration of curves.
@@ 3.1 Space of bounded Expanded degenerations
  1. 1.
    the generic fibre $\mathcal{X}_{\eta}$ is a smooth curve of genus $g\geq 2$,
  2. 2.
@@
Let us denote the relative dualising sheaf by $\omega_{\mathcal{X}/S}$. From here onwards, we assume that $S$ is a neighbourhood of the origin in $\mathbb{A}^{1}:=Spec\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{k}[t]$.
@@
(Gieseker curve/Expanded degeneration/Modification) Let $X_{0}$ be a nodal curve with a single node $x\in X_{0}$, and let $x^{\pm}$ label the two preimages of $x$ in the normalization $\widetilde{X}_{0}$ of $X_{0}$. Let $r$ be a positive integer.
  1. 1.
@@
      for any $i<j$, $R[r]_{i}\cap R[r]_{j}$ consists of a single point $p_{j}$ if $j=i+1$ and empty otherwise.
@@
(Family of Gieseker curves/Expanded degenerations/Modifications:) For every $S$-scheme $T$, a modification of $\mathcal{X}$ relative to $T$ is a commutative diagram
${\mathfrak{X}_{T}}$${\mathcal{X}_{T}:=\mathcal{X}\times_{S}T}$${T}$$\scriptstyle{p_{T}}$$\scriptstyle{\pi_{T}}$ (3.1)
@@
  3. 3.
    over each closed point $t\in T$ which maps to $\eta_{0}\in S$, we have $(\mathfrak{X}_{T})_{t}\cong X_{r}$ for some integer $r$ and the horizontal morphism restricts to the morphism $X_{r}\to X_{0}$ which contracts the $\mathbb{P}^{1}$’s in $X_{r}$ to the node $x\in X_{0}$.
@@
${\mathfrak{X}_{T}}$${\mathfrak{X}^{\prime}_{T}}$${\mathcal{X}_{T}}$$\scriptstyle{\pi_{T}}$$\scriptstyle{\sigma_{T}}$$\scriptstyle{\pi^{\prime}_{T}}$ (3.2)
3.1.2 Space of expanded degenerations
As before we will choose an uniformizer on $S$ so that $S$ is identified with a neighborhood of zero in $\mathbb{A}^{1}$. For any positive integer $n$, we set
$S[n]:=S\times_{\mathbb{A}^{1}}\mathbb{A}^{n+1}$ (3.3)

### v3

@@ Abstract
We construct a flat degeneration of the derived moduli stack of Higgs bundles on smooth projective curves, using Jun Li’s stack of bounded expanded degenerations. The degeneration carries a natural relative zero shifted logarithmic symplectic form over the base, which extends the classical Hitchin symplectic structure from the generic fiber. The induced Hitchin map is shown to be complete (in the valuative sense) and flat, without requiring coprimeness of rank and degree. Finally, we extend the construction globally: the derived moduli stack of Higgs bundles over the universal semi-stable curve of genus greater than or equal to two carries a relative zero shifted log-symplectic form over the moduli stack of stable curves.
1. Introduction
@@
It is well known that moduli stacks of vector bundles on nodal curves are not complete. Often, their completions involve adding torsion-free sheaves as points of the boundary. Because we want to use the mapping space techniques of shifted symplectic geometry, we adopt the alternative approach of completing the moduli by adding boundary points, which are bundles on a bubbling node. This approach goes back to the classical works of David Gieseker [12] and Jun Li [21]. We will implement the approach in our setting by utilizing Jun Li’s stack of expanded degenerations. This allows us to use only vector bundles instead of coherent torsion-free sheaves and apply techniques of shifted symplectic geometry [29], obtaining a relative symplectic structure on a complete moduli stack. We extend the Hitchin map to Higgs bundles on the bubbled curves parametrized by the stack of expanded degenerations and then use the completeness of this moduli stack to prove that our extended Hitchin map is complete. The derived moduli stack of Higgs bundles on a smooth curve $C$ has virtual dimension
$\chi_{\mathrm{Dol}}(C,\mathcal{E}\mathsf{nd}(E,\phi))=n^{2}\deg(c_{1}(T_{C}))=2n^{2}(g-1)$
@@
where $g$ is the genus of $C$, while the derived moduli stack of vector bundles (or any other Lagrangian in the derived moduli stack of the Higgs bundle) has a virtual dimension
$\chi(Y,\mathcal{E}\mathsf{nd}(E))=\int_{C}\text{ch}(\mathcal{E}\mathsf{nd}(E))\text{td}(C)=n^{2}(g-1).$
@@
These moduli stacks are homotopically locally of finite presentation (see [43, §2.2.6.3] and [38]), and so exhibit the ”hidden smoothness” envisioned in the derived deformation theory. Given this, it is natural to study other geometric properties of these moduli. This article investigates the relative 0-shifted symplectic structures in the total space of a degeneration of $C$ to a nodal curve.
@@
After writing this article, we realized that there are many relationships with the articles [4], [3] and significant overlap with a forthcoming article [11] of Ron Donagi and Andres Fernandez Herrero. In their work, they give a construction of a good moduli space for the semistable locus and semistable reduction relative to the Hitchin fibration (so the moduli space is proper over the Hitchin base in families). This uses the “infinite dimensional GIT” picture developed with Dan Halpern-Leistner. They also showed the flatness of the Hitchin morphism and that things are syntomic (as classical stacks), including the symplectic leaves. They are pursuing the story with punctures, thinking about the log Poisson reduction picture for the relative log cotangent stack of framed Gieseker vector bundles.
@@
To start with, we fix a family of projective curves $\mathcal{X}$ on the spectrum of a discrete valuation ring $S$ so that the generic fiber is smooth, the closed fiber is an irreducible nodal curve with a single node, and the total space $\mathcal{X}$ is smooth. We also fix a rank $n$ and degree $d$ for the vector bundles in our moduli problem. In Section §2, we recall the construction of the stack $\mathfrak{M}$ of bounded expanded degenerations (bounded by the integer $n$) of the family of curves $\mathcal{X}/S$ following [21] and [44]. One of the main results in this subsection is the lemma 3.12, which we use to prove the following proposition.
@@
Proposition 3.15 The morphism $\mathfrak{M}\longrightarrow S$ is a log-smooth map, and the relative log-cotangent complex $\mathbb{L}^{log}_{\mathfrak{M}/S}=0$.
@@
In subsection §2.3, we discuss shifted log-symplectic structures on quasi-smooth derived Artin stacks (Definition 3.4.0(a)) equipped with a locally free log structure. In subsection §2.4, we recall the definition of relative shifted symplectic forms for a quasi-smooth morphism of derived Artin stacks. We define relative shifted log-symplectic forms for certain logarithmic morphisms of derived Artin stacks equipped with locally-free log structures. In Section §3, we construct a logarithmic version of the relative Dolbeault moduli stack for the universal expanded family of curves $\mathcal{\mathcal{}}X_{\mathfrak{M}}\rightarrow\mathfrak{M}$. We show that the relative logarithmic Dolbeault moduli stack has a relative 0-shifted log-symplectic form over $S$. Moreover, we show that the relative log-symplectic form is an extension of Hitchin’s symplectic form on the generic fibre of the moduli stack over $S$. This was proved for moduli schemes in [10]. The main results of this section are the following.
@@
Theorem 4.7 $\mathcal{X}_{Dol}$ is $\mathcal{O}$-compact and $\mathcal{O}$-oriented over $S$. Hence, $\mathsf{Map}_{\mathcal{S}}(\mathcal{X}_{Dol},BGL_{n}\times\mathcal{S})$ has a 0-shifted relative symplectic structure over $S$.
@@
Theorem 4.11 The derived Artin stack $\mathsf{Map}_{{}_{\mathfrak{M}}}(\mathcal{X}_{\mathfrak{M},Dol},BGL_{n}\times\mathfrak{M})$ has a natural relative 0-shifted log-symplectic structure over $S$.
@@
In Section §4, we define the Hitchin map on the classical Artin stack of Gieseker-Higgs bundles $M^{cl}_{Gie}$ (see A.1). We prove that the Hitchin map is complete.
Theorem 5.4 The morphism $h|_{M^{cl}_{Gie}}\colon M^{cl}_{Gie}\longrightarrow B$ is complete.
@@
In section §5, we study the reduced global nilpotent cone of $M^{cl}_{Gie}$, which is the reduction of the scheme theoretic fibre over the point $0\in B$. We prove that every irreducible component of the reduced nilpotent cone has an open subset (denoted by $\mathcal{N}ilp^{sm,gen}$), which is an isotropic substack of $M$ (the derived stack of Higgs bundles) with respect to its log-symplectic form. We use this to compute the dimension of the reduced nilpotent cone and to show that the Hitchin map is flat. The main theorem of this section is as follows.
@@
Theorem 7.6 There is a 0-shifted relative log-symplectic form on $\mathcal{M}^{Dol}_{g}$ (relative to the moduli stack of stable curves $\overline{\mathcal{M}_{g}}$).
In the Appendix, we construct the relative classical Artin stack of Gieseker-Higgs bundles and study its local properties. The main results of the appendix are Proposition A.11 and Theorem A.12. In the first lemma, we prove that the stack of Gieseker vector bundles is an almost very good stack. We use this in Lemma A.12 to show that the classical stack of Gieseker-Higgs bundles is an irreducible, local complete intersection.
Proposition A.11 The closed fibre $N^{cl}_{Gie,0}$ is an irreducible, equidimensional, almost very good stack ([39, Definition 2.1.2]) with normal crossing singularities.
Theorem A.12 The stack $M^{cl}_{Gie,0}$ is an irreducible local complete intersection of pure dimension $2\dim N_{Gie,0}+1$.
1.1. Notations and Conventions
@@ 6. $\bullet$
    dgdg, $\mathrm{cdga}$ (respectively $\mathrm{cdga}^{\leq 0}$) are endowed with their natural model structures for which equivalences are quasi-isomorphisms, and fibrations are epimorphisms (respectively, epimorphisms in strictly negative degrees).
@@ 8. $\bullet$
    The $\infty$-categories associated with the model categories $dg,\mathrm{cdga}^{\leq 0},\operatorname{dAff}$ are denoted by $\textbf{dg},\textbf{cdga}^{\leq 0},\textbf{dAff}$.
@@ 10. $\bullet$
    The $\infty$-category of derived stacks over $\mathbb{k}$, for the étale topology, is denoted by dSt. If $X$ is a derived stack, the $\infty$-category of derived stacks over $X$ is denoted by $\textbf{dSt}_{X}$. The classical truncation of a derived stack $X$ is denoted by $t_{0}X$.
@@ 12. $\bullet$
[…]
2. Acknowledgements
T.P. (University of Pennsylvania) was partially supported by NSF FRG grant DMS- 2244978, NSF/BSF grant DMS-2200914, NSF grant DMS-1901876, and Simons Collaboration grant number 347070. O.B. (University of Haifa) and S.D. would like to acknowledge the Binational Science Foundation and NSF/BSF grant DMS-2200914 also known as BSF Grant 2021717 for supporting S.D. as a postdoc. S.D. would like to thank Professor Alek Vainshtein of the University of Haifa, Israel, for the financial support from his Israel Science Foundation grant number 876/20 during a postdoc position at the University of Haifa. S.D. would like to thank Professor Vikraman Balaji for the financial support from his SERB Core Research Grant CRG/2022/000051-G during a postdoc position at Chennai Mathematical Institute. S.D. also thanks IISER Tirupati for their support during a part of this work.
3. Preliminaries
3.1. Space of bounded expanded degenerations
In this subsection, we recall a construction by Jun Li [21] called the stack of expanded degenerations. We will use this stack to construct our degeneration. The construction of Jun Li starts with a one-parameter family of varieties (of any dimension) such that the total space of the family is smooth, the generic fibre is smooth, and the special fibre is a normal crossing divisor in the total space of the family. For our purposes, we only need to consider the expanded degenerations of such a family for curves. We will make use of the standard fact that for any nodal curve, one can always construct a smoothing over the spectrum of a discrete valuation ring, which also has a smooth total space. The precise setting we consider can be explained below.
3.1.1. Degeneration of curves.
@@ 3.1. Space of bounded expanded degenerations
  1. (1)
    generic fiber $\mathcal{X}_{\eta}$ is a smooth curve of genus $g\geq 2$,
  2. (2)
@@
Remark 3.1.
Given any nodal curve, such a family of one-parameter degeneration of curves can always be constructed where the given nodal curve is placed as the closed fiber [1, Theorem B.2 and Corollary B.3, Appendix B]. On the other hand, given any smooth proper family of curves $\mathcal{X}_{K}$ over $\operatorname{Spec}K$, where $K$ is the function field of the discrete valuation ring $S$, one can always extend (possibly after taking a finite cover of $S$) it to a family of stable curves with nodal singularities(possibly with several nodes) [41, Theorem 109.24.3.]. Also note that Jun Li’s construction of the stack of expanded degenerations also works for the case when the closed fiber has more than one node (see [21, page 516]).
Let us denote the relative dualizing sheaf by $\omega_{\mathcal{X}/S}$. From now on, we set $S:=\operatorname{Spec}\penalty 10000\ \mathbb{k}[t]_{(t)}$, which represents an open subset of the affine line $\mathbb{A}^{1}_{\mathbb{k}}$.
@@
(Gieseker curve/Expanded degeneration/Modification) Let $X_{0}$ be a nodal curve with a single node $x\in X_{0}$, and let $x^{\pm}$ label the two preimages of $x$ in normalization $\widetilde{X}_{0}$ of $X_{0}$. Let $r$ be a positive integer.
  1. (1)
@@
      for any $i<j$, $R[r]_{i}\cap R[r]_{j}$ consists of a single point $p_{j}$ if $j=i+1$ and is empty otherwise.
@@
(Family of Gieseker curves / expanded degenerations / modifications:[2, Definition 3.4]) For every $S$-scheme $T$, a modification of $\mathcal{X}$ relative to $T$ is a commutative diagram
(3.1) $\mathfrak{X}_{T}$$\mathcal{X}_{T}:=\mathcal{X}\times_{S}T$$T\scriptstyle{p_{T}}$$\scriptstyle{\pi_{T}}$
@@
  3. (3)
    over each closed point $t\in T$ that maps to $\eta_{0}\in S$, we have $(\mathfrak{X}_{T})_{t}$ is isomorphic to a Gieseker curve $X_{r}$ for some integer $r$ and the horizontal morphism restricts to the morphism $X_{r}\to X_{0}$ that contracts the $\mathbb{P}^{1}$s in $X_{r}$ to the node $x\in X_{0}$.
@@
(3.2) $\mathfrak{X}_{T}$$\mathfrak{X}^{\prime}_{T}$$\mathcal{X}_{T}$$\scriptstyle{\pi_{T}}$$\scriptstyle{\sigma_{T}}$$\scriptstyle{\pi^{\prime}_{T}}$
Remark 3.5.
The definition of a modification or Gieseker curve is stated slightly differently from the definition given in Gieseker’s original paper (see [12, (i) and (iii) of the first paragraph of §4]. The definition given in this paper is the same as the Gieseker’s definition (see [25, (ii) of Definition 7; Lemma. on page 200]). For the deformation theory of Gieseker curves (i.e., the description of such families of modifications over Artin rings), we refer to [12, Proposition 4.5. ], and [25, Local Theory, Page 197].
3.1.2. Space of expanded degenerations
As before, we will choose an uniformizer in $S$ at the closed point and if necessary replace it by $S:=\operatorname{Spec}\penalty 10000\ \mathbb{k}[t]_{(t)}$, which represents an open neibourhood of the origin of the affine line $\mathbb{A}^{1}_{\mathbb{k}}$. For any positive integer $n$, we set
(3.3) $S[n]:=S\times_{\mathbb{A}^{1}}\mathbb{A}^{n+1}$